# Mixtures of Neural Network Experts with Application to Phytoplankton Flow Cytometry Data

### v1

@@ 3.1 Estimation
Following Hyun et al., (2023), we estimate model parameters by maximizing a constrained pseudolikelihood. Phytoplankton flow cytometry data naturally have a class imbalance due to the nature of small, lightweight cells to dominate in number. Larger organisms are less numerous but still account for a significant portion of the biomass in their habitat (Marañón,, 2015). Mixture models tend to overfit to the class which dominates in number (Xu and Jordan,, 1996), so we give additional importance to less numerous but more massive cells by exponentiating the likelihood contribution of each particle by its biomass. This results in the pseudolikelihood:
@@
Here, we use $\alpha$ to denote the collection of coefficients $\{\alpha_{0,k}\}_{k},\{\boldsymbol{\alpha}_{k}\}_{k}$, and $\beta$ to denote $\{\boldsymbol{\beta}_{0,k}\}_{k},\{\boldsymbol{\beta}_{k}\}_{k}$. We also write $\pi_{k,t}(\boldsymbol{x}_{t})$ and $\boldsymbol{\mu}_{k,t}(\boldsymbol{x}_{t})$ to emphasize their dependence on the covariates. Pseudolikelihoods are commonly utilized in informative sampling scenarios, where the likelihood contributions are exponentiated by the observations’ inverse probabilities of selection. The resulting psuedolikelihood converges to the true population-generating distribution under certain regularity conditions (Savitsky and Toth,, 2016). In our dataset, particles with large biomass tend to be less numerous and therefore less likely to be observed, so it is reasonable to assume that a particle’s selection probability is approximately proportional to the inverse of its biomass.
@@
In addition, Hyun et al., (2023) also impose two model constraints to encourage smoothness and penalize overly complex models, a technique we also utilize. First, since the optical properties of any particular species is limited in range due to biological constraints, we constrain the cluster means $\boldsymbol{\mu}_{k,t}$ to vary over time by up to a fixed amount $r$, called the maximum deviation. This constraint can be expressed as
$||\boldsymbol{\mu}_{k,t}-\boldsymbol{\beta}_{0,k}||_{2}=||\boldsymbol{\beta}_{k}^{\prime}\sigma(\boldsymbol{h}_{t})||_{2}\leq r,$
@@
for cluster $k=1,\ldots,K$ and time $t=1,\ldots,T$. Another benefit of this constraint is that a properly chosen value of $r$ often prevents the switching of cluster labels over time. Second, to guard against selecting an overly complex regression model, lasso penalties (Tibshirani,, 2018) are imposed on all regression coefficients. After rescaling the pseudolikelihood, we arrive at the optimization problem
$\begin{split}\min_{\alpha,\beta,\Sigma}-\frac{1}{N}\mathcal{L}(\alpha,\beta,\Sigma;\{\boldsymbol{y}_{i}^{(t)}\}_{i,t},\{\boldsymbol{x_{t}}\}_{t})+\lambda_{\alpha}\sum_{k=1}^{K}||\boldsymbol{\alpha}_{k}||_{1}+\lambda_{\beta}\sum_{k=1}^{K}||\boldsymbol{\beta}_{k}||_{1}\\ \text{subject to }||\boldsymbol{\beta}_{k}^{\prime}\sigma(\boldsymbol{h}_{t})||_{2}\leq r\ \forall k=1,\ldots,K,\ \forall t=1,\ldots,T,\end{split}$ (5)
@@
the solution of which provides parameter estimates. We solve (5) using the penalized EM algorithm described in Hyun et al., (2023) and use cross-validation to select the optimal hyperparameters $\lambda_{\alpha}$ and $\lambda_{\beta}$.
@@
Lastly, since flow cytometry datasets have very large sample sizes $n_{t}$, we follow the approach taken by Hyun et al., (2023) to reduce computational demand by aggregating particles into a fixed grid of $D$ bins in each cytogram dimension. The new binned cytogram at time $t$ consists of observations $\left\{\boldsymbol{y}_{b}^{(t)}:b=1,\ldots,D^{d}\right\}$, where $\boldsymbol{y}_{b}^{(t)}$ is the center location of bin $b$. We associate with bin $b$ a new weight $c_{b}^{(t)}$ equal to the total biomass of particles within the bin, and use $\boldsymbol{y}_{b}^{(t)}$ and $c_{b}^{(t)}$ in place of $\boldsymbol{y}_{i}^{(t)}$ and $c_{i}^{(t)}$ in (5).
@@ 4.1 Simulation design
In this section, we generate pseudo-synthetic data to test the numerical properties of our model. For each signal size $\Delta$ and at each time point $t=1,\cdots,296$, we generate $n_{t}=1,000$ particles $\boldsymbol{y}_{i}^{(t)}$ from a mixture of two one-dimensional Gaussian distributions. Specifically, the latent membership $z_{i}^{(t)}$ is first drawn from $\mathrm{Categorical}(\pi_{1,t},\pi_{2,t})$, and conditional on $z_{i}^{(t)}$, the particle’s value $\boldsymbol{y}_{i}^{(t)}\mid z_{i}^{(t)}=k$ is drawn from $\mathcal{N}\left(\mu_{1,t},0.2^{2}\right)$ if $k=1$, or $\mathcal{N}\left(\mu^{\Delta}_{2,t},0.2^{2}\right)$ if $k=2$. Cluster 1’s mean $\mu_{1,t}$ changes over time $t$ as a function of the two principal components $\boldsymbol{\Psi}_{\cdot,1}$, and $\boldsymbol{\Psi}_{\cdot,4}$, taken from the real data’s principal component matrix $\boldsymbol{\Psi}$ in (1). In contrast, cluster 2’s mean $\mu_{2,t}^{\Delta}$ remains constant over time, so $\mu_{2,t}^{\Delta}=\mu_{2}^{\Delta}$ for all $t$. Also, cluster 1’s cluster probability $\pi_{1,t}$ changes over time $t$ as a function of the two principal components $\boldsymbol{\Psi}_{\cdot,1}$, and $\boldsymbol{\Psi}_{\cdot,2}$, and $\pi_{2,t}=1-\pi_{1,t}$ by construction.
@@
For each configuration, we keep the first mean $\{\mu_{1,t}\}_{t=1}^{296}$ the same and vary the level of the second cluster mean $\mu_{2}^{\Delta}$ over signal sizes $\Delta$ which we define as the size of the gap between the two cluster means (averaged over time). That is, $\mu_{2}^{0}$ with the smallest signal size $\Delta=0$ has a zero gap between the average of the first cluster’s mean $\frac{1}{T}\sum_{t=1}^{T}\mu_{1,t}$ and $\mu_{2}^{\Delta}$. We consider equally spaced signal sizes between zero and the largest value $\Delta=0.95$, the signal size at which the two clusters’ ±$\pm 2$ standard deviation bands around the mean cease to overlap.
@@
        Interaction: $\mu_{1,t}=1.05707+0.015\boldsymbol{\Psi}_{t,1}+0.02\boldsymbol{\Psi}_{t,4}+0.003\boldsymbol{\Psi}_{t,1}\boldsymbol{\Psi}_{t,4}$
@@
        Logistic: $\text{logit}(\pi_{1,t})=-2+4.5(1+\exp(\boldsymbol{\Psi}_{t,1}))^{-1}+0.1\boldsymbol{\Psi}_{t,2}$.
@@
Next, for evaluating predictive out-of-sample accuracy, we use the log-pseudolikelihood in (4), calculated on an independently drawn dataset generated using the same parameters as the training dataset. For all scenarios, we use $5$-fold cross-validation to select tuning parameters, and for each model fit, we run the EM algorithm $30$ times. More specifically, we use a cross-validation scheme designed to prevent the data in the train and test folds from being too close to each other. Rather than using random splits of the time domain $\{1,\cdots,T\}$ for cross-validation folds, we partition it into blocks of 20 consecutive time intervals and then distribute the blocks among the folds. Specifically, for $K=5$ folds, the first fold is assigned blocks $1,6,11,\cdots$ and contain time points $\{1,\cdots,20\}\cup\{121,\cdots,140\}\cup\cdots$, the second fold is assigned the blocks $2,7,12,\cdots$, and so forth.
@@ 4.2 Simulation results
Figure 2 shows an example of the estimated linear model (left panel) and nonlinear model (right panel), both using data generated from a linear probability function and an interaction mean function. In the true model, the variation in the first cluster mean $\{\mu_{1,t}\}_{t=1}^{296}$ over time is largely due to $\boldsymbol{\Psi}_{\cdot,4}$, which is highly correlated with sunlight and thus oscillates regularly with a daily periodicity. The interaction in the linear effect of $\boldsymbol{\Psi}_{\cdot,4}$ with $\boldsymbol{\Psi}_{\cdot,1}$ causes the true amplitude of the oscillation to become smaller at high latitudes. However, the estimated linear model fails to capture this and the amplitude of the estimated oscillation in ${\mu}_{1,t}$ does not shrink enough in high latitudes. It also noticeably underestimates the first cluster mean $\mu_{1,t}$ when $t\in[40,110]$, as well as the second cluster mean $\mu_{2,t}$ across all time points. In contrast, the nonlinear model more accurately captures the change in amplitude, with its estimated means closely matching the ground truth at all time points. Figure 3 shows the aggregate summary of the simulation results. The y-axis values in each panel is the negative log-pseudolikelihood (NLPL) of an estimated model after subtracting the NLPL model of the oracle, or true generating model. It is clear that the nonlinear model almost uniformly outperforms the linear model in terms of the out-of-sample negative log-pseudolikelihood, which demonstrates a better fit to the data. (The nonlinear model even outperforms for linear data sometimes, which may be due to large sample sizes with relatively little noise, which makes training and test datasets very similar.)
@@ 5 Application to SeaFlow cruise data
Figure 4: Cytogram in two dimensions, diameter and chlorophyll, at time $t=33$ (2017-07-02 08:00-09:00), with model estimates overlaid in red (marking the important clusters) or grey (the rest). The data in the background is binned, with the intensity of blue hue proportional to the total biomass in each bin. The solid points mark the cluster means, and the size of the points are proportional to the estimated cluster probabilities. The ellipses represent 2-dimensional projections of symmetric cluster-conditional 95% probability sets of each cluster.
The models’ estimates at the time point $t=33$ are shown in Figure 4, where we focus on two out of the three cytogram dimensions, and overlay the estimated model with the binned data. The resulting clusters from each model have subsequently been expert-identified as Prochlorococcus (Pro), Synechococcus (Syn), and two distinct PicoEukaryote (Pico1 and Pico2), as well as other populations. It is visually clear that the two models identify the major populations (Pro, Syn, Pico1 and Pico2 in Figure 4) similarly.
@@
      Evaluate the unpenalized negative log-pseudolikelihood $-\mathcal{L}\left(\hat{\alpha},\hat{\beta},\hat{\Sigma};\{\boldsymbol{y}_{i}^{(t)}\}_{i,t\in\mathcal{T}^{(j)}},\{\boldsymbol{x}_{t}\}_{t\in\mathcal{T}^{(j)}}\right)$.
@@
    Let $-\frac{1}{5}\sum_{j=1}^{5}\mathcal{L}\left(\hat{\alpha},\hat{\beta},\hat{\Sigma};\{\boldsymbol{y}_{i}^{(t)}\}_{i,t\in\mathcal{T}^{(j)}},\{\boldsymbol{x}_{t}\}_{t\in\mathcal{T}^{(j)}}\right)$ be the final estimate of the out-of-sample model fit.
The linear model had an out-of-sample negative log-pseudolikelihood of $3.446$ while the nonlinear model’s was $3.439$, showing similar generalization performance. In addition, the relative abundance predictions of the two models are very similar, as Figure 5 shows. However, there are notable differences between the two models’ cluster mean predictions. The linear model predicts a dramatic decrease in Prochlorococcus diameter between July 2nd and July 6th, while the nonlinear model predictions remain relatively stable, showing a diel oscillation at a similar overall level. Also, the nonlinear model better captures the amplitude of the expected diel oscillation in chlorophyll of Prochlorococcus throughout all time points, matching its expected behavior from the corresponding oscillation in cell size. Overall, while the models display similar clustering results, the nonlinear model estimates more flexible cluster mean movement over time that fits scientific intuition.
Figure 5: Mean and probability estimates over time for the Prochlorococcus cluster. In the left panel, the three dimensions of the cluster mean are plotted together; diameter, chlorophyll, and phycoerythrin estimates are represented by the red, blue, and yellow lines, respectively. In the right panel, relative abundance predictions are displayed. The nonlinear model predictions are represented by the solid lines and the linear model predictions by dot-dashed lines. The vertical dotted lines represent the boundaries of the North Pacific Transition Zone.
@@
$\hat{\boldsymbol{\mu}}_{k}(\psi_{1},\cdots,\psi_{q})=\boldsymbol{\hat{\beta}}_{0,k}+\boldsymbol{\hat{\beta}}_{k}^{\prime}\sigma(\boldsymbol{W}^{\prime}\left(\begin{smallmatrix}1\\ \psi_{1}\\ \vdots\\ \psi_{q}\end{smallmatrix}\right)),$
@@
Figure 6: Partial responses of Prochlorococcus diameter, the first PicoEukaryote cluster’s log diameter, and the second PicoEukaryote cluster’s relative abundance as functions of the first principal component. The vertical dotted lines represent the boundaries of the North Pacific Transition Zone.
@@
We see in Figure 6 a few representative partial response plots in which certain phytoplankton properties are displayed as functions of PC1. As PC1 increases, the cruise moves northward closer to the subarctic region. We see in the left panel of Figure 6 that both models predict Prochlorococcus diameter to be larger monotonically as one moves to higher latitudes, if one were to hold as constant ocean water conditions that are not associated with latitude (PC1). This strongly suggests that the average Prochlorococcus microbe is larger in higher latitudes. However, the linear model predicts only a flatter and linear change, by design. The nonlinear model is capable of predicting a more flexible change. It predicts Prochlorococcus diameter to change (1) much slower in the south-most and north-most regions along the cruise track, and (2) much faster in the transition zone, especially within the NPTZ. This is consistent with external studies (Thompson et al.,, 2021) that different subtypes of Prochlorococcus are dominant in each of the two gyres – Subpolar and Subtropical – and our model estimates further suggest that those subtypes have a clear size difference.
For PicoEukaryote, both models predict that the cell diameter decreases monotonically as one goes from south to north along the cruise trajectory. As with Prochlorococcus, the cell diameter levels off to become flat for high latitudes, indicating stable populations of PicoEukaryote at high latitudes. In fact, when PC1 approaches ±$\pm 8$, both PicoEukaryote and Prochlorococcus clusters’ cell diameters are predicted to flatten, reflecting the existence of maximum and minimum sizes of plankton cells. The linear model predictions, however, do not level off, and any prediction outside of the observed latitude range is sure to be too extreme to be scientifically useful.
Finally, the second PicoEukaryote cluster (Pico2) is predicted to have a maximum relative abundance at a value of PC1 which corresponds to latitude 38.8 N. At more northern locations than this, the relative abundance is predicted to decrease. Such results highlight the possibility that nonlinear mixture-of-experts models, such as our proposed model, can be used to learn globally and locally optimal ocean conditions where certain phytoplankton species are most relatively abundant.
Figure 7: The top panel displays partial response curves of certain model parameters as functions of the second principal component, conditional upon several different values of the first principal component. The bottom panel displays the trajectory of the Gradients 2 cruise, with horizontal lines marking the values of PC1 used in the top panel as well as the NPTZ boundaries.
@@
over a fine grid of values of $x$ in the range of the second principal component $\boldsymbol{\Psi}_{\cdot,2}$. Figure 7 highlights the conditional partial responses of the cluster probability $\hat{\pi}_{k,t}$ (in the left panel) or one dimension of the cluster mean $\hat{\boldsymbol{\mu}}_{k,t}$ (in the center and right panel), fixing the values of PC1 at three values ($-6$, $-1$, and $4$) which correspond to three locations clearly south of, inside of, and north of the North Pacific Transition Zone.
First, for cluster means, whenever the three lines from the same model are not parallel and equidistant, this indicates an interaction between PC1 and PC2. We see this most clearly in PicoEukaryote diameter in the right panel. At a high value of PC1 (4, corresponding to north of the NPTZ), the partial response curve is nearly linear. However, at lower values of PC1 (inside or south of the NPTZ), the partial response to PC2 clearly becomes nonlinear. In the center panel showing the phycoerythrin fluorescence of Synechococcus, the partial response curve changes from convex for lower PC1 values to concave for higher PC1 values.
Next, we examine the relative abundance prediction of Prochlorococcus (left panel of Figure 7). First, the linear model’s partial response curve to PC2 follows a similar logistic shape at all specified values of PC1. On the other hand, the nonlinear model curve clearly changes in shape, especially from PC1 = $-1$ to PC1 = $-6$. This suggests an interaction effect between PC1 and PC2 which allows the response of Prochlorococcus relative abundance to PC2 to be much more flexible depending on the latitude, of which PC1 serves as a proxy. Furthermore, the nonlinear model has a shallower slope overall, thus estimating Prochlorococcus abundance to be more stable in response to physical oceanographic attributes (PC2) in southern waters than the linear model does.

### v2

@@ 3.1 Estimation
Following [24], we estimate model parameters using the frequentist approach of maximizing a constrained pseudolikelihood. Phytoplankton flow cytometry data naturally have a class imbalance due to the nature of small, lightweight cells to dominate in number. Larger organisms are less numerous but still account for a significant portion of the biomass in their habitat [30]. Mixture models tend to overfit to the class which dominates in number and therefore underestimate the abundance of rare classes [47], so we give additional importance to less numerous but more massive cells by exponentiating the likelihood contribution of each particle by its biomass. This results in the pseudolikelihood:
@@
Here, we use $\alpha$ to denote the collection of coefficients $\{\alpha_{0,k}\}_{k},\{\boldsymbol{\alpha}_{k}\}_{k}$, and $\beta$ to denote $\{\boldsymbol{\beta}_{0,k}\}_{k},\{\boldsymbol{\beta}_{k}\}_{k}$. We also write $\pi_{k,t}(\boldsymbol{x}_{t})$ and $\boldsymbol{\mu}_{k,t}(\boldsymbol{x}_{t})$ to emphasize their dependence on the covariates. Pseudolikelihoods are commonly utilized in informative sampling scenarios, where the likelihood contributions are exponentiated by the observations’ inverse probabilities of selection. The resulting psuedolikelihood converges to the true population-generating distribution under certain regularity conditions [37]. In our dataset, particles with large biomass tend to be less numerous and therefore less likely to be observed, so it is reasonable to assume that a particle’s selection probability is approximately proportional to the inverse of its biomass. Supplementary Figure S13 in Supplement G shows that the biomass weights are effective in improving the estimated distribution of biomass in the SeaFlow application.
@@
In addition, [24] also impose two model constraints to encourage smoothness and penalize overly complex models, a technique we also utilize. First, since the optical properties of any particular species is limited in range due to biological constraints, we constrain the cluster means $\boldsymbol{\mu}_{k,t}$ to vary over time by up to a fixed amount $r$, called the maximum deviation. This constraint can be expressed as
[…]
@@
for cluster $k=1,\ldots,K$ and time $t=1,\ldots,T$. Another benefit of this constraint is that a properly chosen value of $r$ often prevents the switching of cluster labels over time. Since (4) restricts the cluster means to a hypersphere, the data must be scaled so that each response dimension has roughly equal range. Second, to guard against selecting an overly complex regression model, lasso penalties [42] are imposed on all regression coefficients. Now, maximizing the constrained and penalized pseudolikelihood can also be formulated as the following optimization problem:
[…]
@@
the solution of which provides parameter estimates. We solve (5) using the penalized EM algorithm described in [24] and use cross-validation to select the optimal hyperparameters $\lambda_{\alpha}$ and $\lambda_{\beta}$. To avoid data leakage between train and validation folds, we learn the PCA transformation from only the training data within each fold, though $q=9$ is fixed a priori as in Section 3.
@@
Lastly, since flow cytometry datasets have very large sample sizes $n_{t}$, we follow the approach taken by [24] to reduce computational demand by aggregating particles into a fixed grid of $D$ bins in each cytogram dimension. The new binned cytogram at time $t$ consists of observations $\left\{\boldsymbol{y}_{b}^{(t)}:b=1,\ldots,D^{d}\right\}$, where $\boldsymbol{y}_{b}^{(t)}$ is the center location of bin $b$. We associate with bin $b$ a new weight $c_{b}^{(t)}$ equal to the total biomass of particles within the bin, and use $\boldsymbol{y}_{b}^{(t)}$ and $c_{b}^{(t)}$ in place of $\boldsymbol{y}_{i}^{(t)}$ and $c_{i}^{(t)}$ in (5). The total number of non-empty bins in the SeaFlow dataset is $N=\sum_{t=1}^{T}n_{t}=1,178,236$, which highlights the need for a computationally efficient nonlinear regression framework.
@@ 4.1 Simulation design
In this section, we generate pseudo-synthetic data to test the numerical properties of our model. For each signal size $\Delta$ and at each time point $t=1,\cdots,296$, we generate $n_{t}=1,000$ particles $\boldsymbol{y}_{i}^{(t)}$ from a mixture of two one-dimensional Gaussian distributions. Specifically, the latent membership $z_{i}^{(t)}$ is first drawn from $\mathrm{Categorical}(\pi_{1,t},\pi_{2,t})$, and conditional on $z_{i}^{(t)}$, the particle’s value $\boldsymbol{y}_{i}^{(t)}\mid z_{i}^{(t)}=k$ is drawn from $\mathcal{N}\left(\mu_{1,t},0.2^{2}\right)$ if $k=1$, or $\mathcal{N}\left(\mu^{\Delta}_{2,t},0.2^{2}\right)$ if $k=2$. Cluster 1’s mean $\mu_{1,t}$ changes over time $t$ as a function of the two principal components $\boldsymbol{\Psi}_{\cdot,1}$, and $\boldsymbol{\Psi}_{\cdot,4}$, taken from the real data’s principal component matrix $\boldsymbol{\Psi}$ in 2. In contrast, cluster 2’s mean $\mu_{2,t}^{\Delta}$ remains constant over time, so $\mu_{2,t}^{\Delta}=\mu_{2}^{\Delta}$ for all $t$. Also, cluster 1’s cluster probability $\pi_{1,t}$ changes over time $t$ as a function of the two principal components $\boldsymbol{\Psi}_{\cdot,1}$, and $\boldsymbol{\Psi}_{\cdot,2}$, and $\pi_{2,t}=1-\pi_{1,t}$ by construction.
@@
For each configuration, we keep the first mean $\{\mu_{1,t}\}_{t=1}^{296}$ the same and vary the level of the second cluster mean $\mu_{2}^{\Delta}$ over signal sizes $\Delta$ which we define as the size of the gap between the two cluster means (averaged over time). That is, $\mu_{2}^{0}$ with the smallest signal size $\Delta=0$ has a zero gap between the average of the first cluster’s mean $\frac{1}{T}\sum_{t=1}^{T}\mu_{1,t}$ and $\mu_{2}^{\Delta}$. We consider equally spaced signal sizes between zero and the largest value $\Delta=0.95$, the signal size at which the two clusters’ $\pm 2$ standard deviation bands around the mean cease to overlap.
@@
        Interaction: $\mu_{1,t}=1.05707+0.015\boldsymbol{\Psi}_{t,1}+0.02\boldsymbol{\Psi}_{t,4}-0.003\boldsymbol{\Psi}_{t,1}\boldsymbol{\Psi}_{t,4}$
@@
        Logistic: $\text{logit}(\pi_{1,t})=-2-3.5(1+\exp(\boldsymbol{\Psi}_{t,1}))^{-1}+0.1\boldsymbol{\Psi}_{t,2}$.
@@
Next, for evaluating predictive out-of-sample accuracy, we use the log-pseudolikelihood in (3), calculated on an independently drawn dataset generated using the same parameters as the training dataset. For all scenarios, we use $5$-fold cross-validation to select tuning parameters, and for each model fit, we run the EM algorithm $30$ times. More specifically, we use a cross-validation scheme designed to prevent the data in the train and test folds from being too close to each other. Rather than using random splits of the time domain $\{1,\cdots,T\}$ for cross-validation folds, we partition it into blocks of 20 consecutive time points and then distribute the blocks among the folds. Specifically, for $K=5$ folds, the first fold is assigned blocks $1,6,11,\cdots$ and contain time points $\{1,\cdots,20\}\cup\{121,\cdots,140\}\cup\cdots$, the second fold is assigned the blocks $2,7,12,\cdots$, and so forth.
@@ 4.2 Simulation results
Figure 2 shows an example of the estimated linear model (left panel) and nonlinear model (right panel), both using data generated from a linear probability function and an interaction mean function. In the true model, the variation in the first cluster mean $\{\mu_{1,t}\}_{t=1}^{296}$ over time is largely due to $\boldsymbol{\Psi}_{\cdot,4}$, which is highly correlated with sunlight and thus oscillates regularly with a daily periodicity. The interaction in the linear effect of $\boldsymbol{\Psi}_{\cdot,4}$ with $\boldsymbol{\Psi}_{\cdot,1}$ causes the true amplitude of the oscillation to become smaller at high latitudes. However, the estimated linear model fails to capture this and the amplitude of the estimated oscillation in ${\mu}_{1,t}$ does not shrink enough in high latitudes. It also noticeably underestimates the first cluster mean $\mu_{1,t}$ when $t\in[40,110]$, as well as the second cluster mean $\mu_{2,t}$ across all time points. In contrast, the nonlinear model more accurately captures the change in amplitude, with its estimated means closely matching the ground truth at all time points. Supplement E shows that the nonlinear model results are reproducible across a variety of hidden layer widths (Supplementary Figure S10) and draws of the random weights (Supplementary Figure S11).
Figure 3 shows the aggregate summary of the simulation results. The y-axis values in each panel is the negative log-pseudolikelihood (NLPL) of an estimated model after subtracting the NLPL model of the oracle, or true generating model. It is clear that the nonlinear model almost uniformly outperforms the linear model in terms of the out-of-sample negative log-pseudolikelihood, which demonstrates a better fit to the data. (The nonlinear model even outperforms for linear data sometimes, which may be due to large sample sizes with relatively little noise, which makes training and test datasets very similar.) We also provide a simulation with three-dimensional response data in Supplement D, which highlights the ability of our proposed model to recover multivariate Gaussian mixtures under heavy cluster overlap. The model fit to the data is shown in Supplementary Figure S8, and a comparison of the true and estimated parameters is provided in Supplementary Figure S9.
@@ 5 Application to SeaFlow cruise data
Figure 4: Cytogram in two dimensions, diameter and chlorophyll, at time $t=33$ (2017-07-02 08:00-09:$0$), with model estimates overlaid in red (marking the important clusters) or grey (the rest). The data in the background is binned, with the intensity of blue hue proportional to the total biomass in each bin. The solid points mark the cluster means, and the size of the points are proportional to the estimated cluster probabilities. The ellipses represent 2-dimensional projections of symmetric cluster-conditional 95% probability sets of each cluster.
The models’ estimates at the time point $t=33$ are shown in Figure 4, where we focus on two out of the three cytogram dimensions, and overlay the estimated model with the binned data. The resulting clusters from each model have subsequently been expert-identified as Prochlorococcus (Pro), Synechococcus (Syn), and two distinct PicoEukaryote (Pico1 and Pico2), as well as other populations. It is visually clear that the two models identify the major populations (Pro, Syn, Pico1 and Pico2 in Figure 4) similarly. The nonlinear model predicts the PicoEukaryote 1 cluster to be bigger, which captures more of the biomass between the Synechococcus and PicoEukaryote 2 clusters.
@@
      Evaluate the unpenalized negative log-pseudolikelihood $\mathcal{L}\left(\hat{\alpha},\hat{\beta},\hat{\Sigma};\{\boldsymbol{y}_{i}^{(t)}\}_{i,t\in\mathcal{T}^{(j)}},\{\boldsymbol{x}_{t}\}_{t\in\mathcal{T}^{(j)}}\right)$.
@@
    Let $\frac{1}{5}\sum_{j=1}^{5}\mathcal{L}\left(\hat{\alpha},\hat{\beta},\hat{\Sigma};\{\boldsymbol{y}_{i}^{(t)}\}_{i,t\in\mathcal{T}^{(j)}},\{\boldsymbol{x}_{t}\}_{t\in\mathcal{T}^{(j)}}\right)$ be the final estimate of the out-of-sample model fit.
The linear model had an out-of-sample negative log-pseudolikelihood of $3.486$ while the nonlinear model’s was $3.414$, showing similar generalization performance. In addition, the relative abundance predictions of the two models are very similar, as Figure 5 shows. However, there are notable differences between the two models’ cluster mean predictions. The linear model predicts a dramatic decrease in Prochlorococcus diameter between July 2nd and July 6th, while the nonlinear model predictions remain relatively stable, showing a diel oscillation at a similar overall level. Also, the nonlinear model better captures the amplitude of the expected diel oscillation in chlorophyll of Prochlorococcus throughout all time points, matching its expected behavior from the corresponding oscillation in cell size. Overall, while the models display similar clustering results, the nonlinear model estimates more flexible cluster mean movement over time that fits scientific intuition.
Figure 5: Mean and probability estimates over time for the Prochlorococcus cluster. In the left panel, the three dimensions of the cluster mean are plotted together; diameter, chlorophyll, and phycoerythrin estimates are represented by the red, blue, and yellow lines, respectively. In the right panel, relative abundance predictions are displayed. The nonlinear model predictions are represented by the solid lines and the linear model predictions by dot-dashed lines. The vertical dotted lines represent the boundaries of the North Pacific Transition Zone. Supplementary Figures S3–-S5 in Supplement B present similar plots for Synechococcus and PicoEukaryote clusters 1 and 2, respectively.
@@
$\hat{\boldsymbol{\mu}}_{k}(\psi_{1},\cdots,\psi_{q})=\boldsymbol{\hat{\beta}}_{0,k}+\boldsymbol{\hat{\beta}}_{k}^{\prime}\sigma(\boldsymbol{W}\left(\begin{smallmatrix}1\\ \psi_{1}\\ \vdots\\ \psi_{q}\end{smallmatrix}\right)),$
@@
Figure 6: Partial responses of Prochlorococcus diameter, Synechococcus phycoerythrin fluorescence, and the second PicoEukaryote cluster’s relative abundance as functions of the first principal component. The vertical dotted lines represent the boundaries of the North Pacific Transition Zone. The full set of partial response curves for PC1 is provided in Supplementary Figure S6 (Supplement C).
@@
We see in Figure 6 a few representative partial response plots in which certain phytoplankton properties are displayed as functions of PC1. As PC1 increases, the cruise moves northward closer to the subarctic region. We see in the left panel of Figure 6 that both models predict Prochlorococcus diameter to be larger monotonically as one moves to higher latitudes, if one were to hold as constant ocean water conditions that are not associated with latitude (PC1). This strongly suggests that the average Prochlorococcus microbe is larger in higher latitudes. However, the linear model predicts only a flatter and linear change, by design. The nonlinear model is capable of predicting a more flexible change. It predicts Prochlorococcus diameter to change (1) much slower in the south-most and north-most regions along the cruise track, and (2) much faster in the transition zone, especially within the NPTZ. This is consistent with external studies [41] that different subtypes of Prochlorococcus are dominant in each of the two gyres – Subpolar and Subtropical – and our model estimates further suggest that those subtypes have a clear size difference.
For Synechococcus, both models predict that phycoerythrin fluorescence increases monotonically as one goes from south to north along the cruise trajectory. As with Prochlorococcus diameter, the nonlinear model curve flattens at extreme latitudes, indicating stable biological function of Synechococcus cells. The linear model predictions, however, do not level off, and any prediction outside of the observed latitude range is sure to be too extreme to be scientifically useful. This finding is especially important since the presence of phycoerythrin in a cell uniquely identifies it as Synechococcus, since other phytoplankton contain little to none of this pigment.
Finally, the second PicoEukaryote cluster (Pico2) is predicted to have a maximum relative abundance at a value of PC1 which corresponds to latitude 38.8 N. At more northern locations than this, the relative abundance is predicted to decrease. Such results highlight the possibility that nonlinear mixture-of-experts models, such as our proposed model, can be used to learn globally and locally optimal ocean conditions where certain phytoplankton species are most relatively abundant. In Supplement F, we refit the nonlinear model for different numbers of principal components; Supplementary Figure S12 shows that the partial response curves remain relatively stable. We conclude that model interpretation is robust to the number of principal components input to the neural network.
Figure 7: The top panel displays partial response curves of certain model parameters as functions of the second principal component, conditional upon several different values of the first principal component. The bottom panel displays the trajectory of the Gradients 2 cruise, with horizontal lines marking the values of PC1 used in the top panel as well as the NPTZ boundaries. The full set of partial response curves for PC2, conditional on PC1, is provided in Supplementary Figure S7 (Supplement C).
@@
over a fine grid of values of $x$ in the range of the second principal component $\boldsymbol{\Psi}_{\cdot,2}$. Figure 7 highlights the conditional partial responses of the cluster probability $\hat{\pi}_{k,t}$ (in the left panel) or one dimension of the cluster mean $\hat{\boldsymbol{\mu}}_{k,t}$ (in the center and right panel), fixing the values of PC1 at three values (6-6, 1-1, and $4$) which correspond to three locations clearly south of, inside of, and north of the North Pacific Transition Zone.
First, for cluster means, whenever the three lines from the same model are not parallel and equidistant, this indicates an interaction between PC1 and PC2. We see this most clearly in Prochlorococcus chlorophyll fluorescence in the right panel. At a low value of PC1 (-6, corresponding to south of the NPTZ), the partial response curve is relatively flat, convex, and quadratic. At PC1 = -1 (inside the NPTZ), fluorescence increases monotonically at a faster rate, and at PC1 = 4 (north of the NPTZ) the curve becomes concave, meaning fluorescence increases at a decreasing rate. Since the linear model cannot estimate variable rates of change, it averages the increasing and decreasing effects, so fluorescence is estimated to be constant in all three regions. In the center panel showing the phycoerythrin fluorescence of Synechococcus, the partial response curve shows a much faster rate of decrease within the transition zone than outside it.
Next, we examine the relative abundance prediction of Prochlorococcus (left panel of Figure 7). First, the linear model’s partial response curve to PC2 follows a similar logistic shape at all specified values of PC1. On the other hand, the nonlinear model curve clearly changes in shape, especially from PC1 = 1-1 to PC1 = 6-6. This suggests an interaction effect between PC1 and PC2 which allows the response of Prochlorococcus relative abundance to PC2 to be much more flexible depending on the latitude, of which PC1 serves as a proxy. Furthermore, the nonlinear model has a shallower slope overall, thus estimating Prochlorococcus abundance to be more stable in response to physical oceanographic attributes (PC2) in southern waters than the linear model does.